OpenStudy (anonymous):

what will be the closed formula of the series from k = 0 to n of kr^k?

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}kr^k\]?

OpenStudy (anonymous):

think the gimmick is this looks like the derivative of \[r^k\] but you are off by one in the exponent

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so first write \[\sum_{k=0}^{\infty}r^k=\frac{1}{1-r}\]

OpenStudy (anonymous):

then take the derivative to get \[\sum kr^{k-1}=\frac{1}{(1-r)^2}\]

OpenStudy (anonymous):

then multiply by r to get \[\sum kr^k =\frac{r}{(1-r)^2}\]

OpenStudy (anonymous):

as i recall that is the gimmick used here

OpenStudy (anonymous):

you can basically ignore the problem about starting at k =0 or k = 1 because in this case k = 0 gives 0, so don't worry about that part.

OpenStudy (anonymous):

why did we writ it as 1/(1-r)?

OpenStudy (anonymous):

you mean \[\sum r^k=\frac{1}{1-r}\]?

OpenStudy (anonymous):

geometric series

OpenStudy (anonymous):

i was assuming that this was well known

OpenStudy (anonymous):

what alchemista said

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Geometric_series

OpenStudy (anonymous):

if we have to prove this then we have lots of work to do, but the question led me to believe that this was basically taken for granted

OpenStudy (anonymous):

The proof of the closed form for a geometric series is in that wiki article.

OpenStudy (anonymous):

Thanks y'all!